# MIBench: A Comprehensive Benchmark for Model Inversion Attack and Defense

### v2

@@ Abstract
Model Inversion (MI) attacks aim at leveraging the output information of target models to reconstruct privacy-sensitive training data, raising widespread concerns on privacy threats of Deep Neural Networks (DNNs). Unfortunately, in tandem with the rapid evolution of MI attacks, the lack of a comprehensive, aligned, and reliable benchmark has emerged as a formidable challenge. This deficiency leads to inadequate comparisons between different attack methods and inconsistent experimental setups. In this paper, we introduce the first practical benchmark for model inversion attacks and defenses to address this critical gap, which is named MIBench. This benchmark serves as an extensible and reproducible modular-based toolbox and currently integrates a total of 16 state-of-the-art attack and defense methods. Moreover, we furnish a suite of assessment tools encompassing 9 commonly used evaluation protocols to facilitate standardized and fair evaluation and analysis. Capitalizing on this foundation, we conduct extensive experiments from multiple perspectives to holistically compare and analyze the performance of various methods across different scenarios, which overcomes the misalignment issues and discrepancy prevalent in previous works. Based on the collected attack methods and defense strategies, we analyze the impact of target resolution, defense robustness, model predictive power, model architectures, transferability and loss function. Our hope is that this MIBench could provide a unified, practical and extensible toolbox and is widely utilized by researchers in the field to rigorously test and compare their novel methods, ensuring equitable evaluations and thereby propelling further advancements in the future development.
@@ 1 Introduction
In recent years, Model Inversion (MI) attacks have raised alarms over the potential privacy breaches of sensitive personal information, including the leakage of privacy images in face recognition models (He et al., 2016), sensitive health details in medical data (Wang et al., 2022), financial information such as transaction records and account balances (Ozbayoglu et al., 2020), and personal preferences and social connections in social media data (Feng et al., 2022). In the MI attacks, an attacker aims to infer private training data from the output information of the target model. Fredrikson et al. (2014) proposed the first MI attack against linear regression models to reconstruct sensitive features of genomic data. Subsequent studies (Fredrikson et al., 2015; Song et al., 2017; Yang et al., 2019) have extended MI attacks to more Machine Learning (ML) models. Zhang et al. (2020b) first introduces the GANs as stronger image priors, paving the way for more applications in GAN-based methods (Chen et al., 2021a; Yuan et al., 2022; 2023b; Wang et al., 2021b).
@@
However, some critical challenges are posed owing to the lack of a comprehensive, fair, and reliable benchmark. Specifically, the evaluation of new methods is often confined to comparisons with a narrow selection of prior works, limiting the scope and depth of analysis. For instance, some methods (Nguyen et al., 2023b; Yuan et al., 2023a) exhibit superior performance for lower-resolution images while other methods (Struppek et al., 2022; Qiu et al., 2024) perform better at higher resolutions. However, these studies only evaluate under their predominant resolutions, and thus do not provide a unified and comprehensive comparison. Additionally, the absence of unified experimental protocols results in a fragmented landscape where there is less validity and fairness in the comparative studies. For example, the original GMI (Zhang et al., 2020a) achieves the attack accuracy of 28%, 44%, 46% when attacking three different classifiers trained on the CelebFaces Attributes(CelebA) (Liu et al., 2015), while merely maintaining lower attack accuracy of 21%, 32%, 31% and 21%, 31%, 29% when compared in the KEDMI (Chen et al., 2021b) and PLGMI (Yuan et al., 2023a) under the same experimental setup. Compounding these issues, discrepancies in the evaluation metrics further obscure the reliability of reported conclusions, potentially steering the field towards biased or misleading insights. These shortcomings impede both the accurate measurement of advancements in the MI field and the systematic exploration of its theoretical underpinnings, underscoring the urgent need for a harmonized framework to facilitate robust and transparent research practices.
@@
To alleviate these problems, we establish the first benchmark in MI field, named MIBench. For building an extensible modular-based toolbox, we disassemble the pipeline of MI attacks and defenses into four main modules, each designated for data preprocessing, attack methods, defense strategies and evaluation, hence enhancing the extensibility of this unified framework. The proposed MIBench has encompassed a total of 16 distinct attack and defense methods, coupled with 9 prevalent evaluation protocols to adequately measure the comprehensive performance of individual MI methods. Furthermore, we conduct extensive experiments from diverse perspectives to achieve a thorough appraisal of the competence of existing MI methods, while simultaneously venturing into undiscovered insights to inspire potential avenues for future research. We expect that this reproducible benchmark will facilitate the further development of MI field and bring more innovative explorations in the subsequent study. Our main contributions are as follows:
We build the first comprehensive benchmark in MI field, which serves as an extensible and reproducible modular-based toolbox for researchers. We expect that this will foster the development of more powerful MI attacks by making it easier to evaluate their effectiveness across multiple distinct dimensions.
We implement 16 state-of-the-art attack methods and defense strategies and 9 evaluation protocols currently in our benchmark. We hope that the benchmark can further identify the most successful ideas in defending against rapid development of potential MI attacks.
We conduct extensive experiments to thoroughly assess different MI methods under multiple settings and study the effects of different factors to offer new insights on the MI field. In particular, we validate that stronger model predictive power correlates with an increased likelihood of privacy leakage. Moreover, our analysis reveals that certain defense algorithms also fail when the target model achieves high prediction accuracy.
@@ 2 Related Work
Model Inversion Attacks. In the MI attacks, the malicious adversary aims to reconstruct privacy-sensitive data by leveraging the output prediction confidence of the target classifier and other auxiliary priors. Normally, the attacker requires a public dataset that shares structural similarities with the private dataset but without intersecting classes to pre-train the generator. For example, an open-source face dataset serves as essential public data when targeting a face recognition model. For a typical GAN-based MI attack, attackers attempt to recover private images from the GAN’s latent vectors initialized by Gaussian distribution, given the target image classifier parameterized with weights and the trained generator . The attack process can be formulated as follows:
@@
where is the target class, typically denotes the classification loss, is a hyperparameter, and is the prior knowledge regularization (e.g., the discriminator’s classification loss) used to improve the reality of . By minimizing the above equation, the adversary updates the latent vectors into the optimal results and generate final images through .
@@
Fredrikson et al. (2014) first introduce the concept of MI attacks in the context of genomic privacy. They find that maximizing the posterior probabilities of a linear regression model can reconstruct the original genomic markers. Ensuing works (Fredrikson et al., 2015; Song et al., 2017; Yang et al., 2019) manage to design MI attacks for more kinds of models and private data, but are still limited to attacking simple networks and grayscale images. To enhance the reconstruction performance on complex RGB images, GMI (Zhang et al., 2020a) first propose to incorporate the rich prior knowledge (Fang et al., 2023; Gu et al., 2020; Fang et al., 2024) within the pre-trained Generative Adversarial Networks (GANs) (Goodfellow et al., 2014). Specifically, GMI starts by generating a series of preliminary fake images, and then iteratively optimizes the input latent vectors that are used for generation. Based on GMI, KEDMI (Chen et al., 2021b) refine the discriminator by introducing target labels to recover the distribution of the input latent vectors. VMI (Wang et al., 2021a) utilize variational inference to model MI attacks and adopts KL-divergence as the regularization to better estimate the target distribution. PPA (Struppek et al., 2022) introduce a series of techniques such as initial selection, post-selection, and data argumentation to enhance MI attacks and manages to recover high-resolution images by the pre-trained StyleGAN2 (Karras et al., 2019). LOMMA (Nguyen et al., 2023b) integrate model augmentation and model distillation into MI attacks to tackle the problem of over-fitting. PLGMI (Yuan et al., 2023a) leverage a top- selection technique to generate pseudo labels to further guide the training process of GAN.
Besides, based on whether the parameters and structures of the victim model are fully accessible to the attackers, MI attacks can be further split into white-box attacks and black-box attacks. Note that in black-box settings, the gradients can no longer be computed by the back-propagation process. Thus, Yuan et al. (2022) address this problem by sampling numerous latent vectors from random noise, selecting the ones that can generate correct labels, and updating the latent vectors solely on discriminator loss. Nguyen et al. (2023a) propose to conduct MI attacks on various surrogate models instead of the unfamiliar victim model, transforming the black-box settings into white-box settings.
@@
Model Inversion Defenses. To defend MI attacks, most existing methods can be categorized into two types: model output processing (Yang et al., 2020; Wen et al., 2021; Ye et al., 2022) and robust model training (Gong et al., 2023; Titcombe et al., 2021; Li et al., 2022; Wang et al., 2021c; Peng et al., 2022; Struppek et al., 2023). Model output processing refers to reducing the private information carried in the victim model’s output to promote privacy. Yang et al. (2020) propose to train an autoencoder to purify the output vector by decreasing its degree of dispersion. Wen et al. (2021) apply adversarial noises to the model output and confuse the attackers. Ye et al. (2022) leverage a differential privacy mechanism to divide the output vector into multiple sub-ranges. Robust model training refers to that incorporating the defense strategies during the training process. MID Wang et al. (2021c) penalizes the mutual information between model inputs and outputs in the training loss, thus reducing the redundant information carried in the model output that may be abused by the attackers. However, simply decreasing the dependency between the inputs and outputs also results in model performance degradation. To alleviate this issue and strike a better balance between model utility and user privacy, Bilateral Dependency Optimization (BiDO) (Peng et al., 2022) minimizes the dependency between the inputs and outputs while maximizing the dependency between the latent representations and outputs. Gong et al. (2023) propose to leverage GAN to generate fake public samples to mislead the attackers. Titcombe et al. (2021) defend MI attacks by adding Laplacian noise to intermediate representations. LS (Struppek et al., 2023) finds that label smoothing with negative factors can help privacy preservation. TL (Ho et al., 2024) leverages transfer learning to limit the number of layers encoding sensitive information and thus improves the robustness to MI attacks.
@@ 3.1 Dataset
Considering existing MI attacks primarily focus on reconstructing private facial data from image classifiers, we select 4 widely recognized face datasets as the basic datasets for our benchmark, which include Flickr-Faces-HQ (FFHQ) (Karras et al., 2019), MetFaces (Karras et al., 2020a), FaceScrub (Ng & Winkler, 2014), and CelebFaces Attributes (CelebA) (Liu et al., 2015). Generally, FFHQ and MetFaces are employed as public datasets for pre-training auxiliary priors, whereas FaceScrub and CelebA serve as the target private datasets to attack. Our benchmark facilitates researchers to freely combine public datasets with private ones, thereby enabling customized experimental setups. Extensive evaluation on more non-facial datasets are presented in Sec.C.3
Notably, the target resolutions across different MI attacks are not uniform. The majority of attack methods concentrate on low-resolution images of , while recent attack methods have begun to focus on higher resolutions, such as . Therefore, our benchmark offers 2 versions of low-resolution and higher-resolution for the aforementioned 4 datasets and prepares multiple transformation tools for processing images, freeing researchers from the laborious tasks of data preprocessing. More details regarding the datasets can be found in the Sec. B.2 of the Appendix.
@@ 3.2 Implemented Methods
Our benchmark includes a total of 16 methods, comprising 11 attack methods and 4 defense strategies. With a focus on Generative Adversarial Network (GAN)-based MI attacks, we selectively reproduce methods from recent years that have been published in top-tier conferences or journals in the computer vision or machine learning domains. This criterion ensures the reliability and validity of the implemented methods. Considering the main targets in our benchmark are image classifiers for RGB images, the learning-based MI attacks (Fredrikson et al., 2015; Song et al., 2017; Yang et al., 2019) are not incorporated currently. More detailed information about the implemented methods is stated in Sec. A.2 and Sec. A.3.
@@
| GMI (Zhang et al., 2020a) | White-box | CVPR-2020 | WGAN-GP | |
| KEDMI (Chen et al., 2021b) | White-box | ICCV-2021 | Inversion-specific GAN† | |
| VMI (Wang et al., 2021a) | White-box | NeurIPS-2021 | StyleGAN2 | |
| Mirror∗ (An et al., 2022) | White-box/Black-box | NDSS-2022 | StyleGAN | |
| PPA (Struppek et al., 2022) | White-box | ICML-2022 | StyleGAN2 | |
| PLGMI (Yuan et al., 2023a) | White-box | AAAI-2023 | Conditional GAN | |
| LOMMA (Nguyen et al., 2023b) | White-box | CVPR-2023 | ||
| IF-GMI (Qiu et al., 2024) | White-box | ECCV-2024 | StyleGAN2 | |
| BREPMI (Kahla et al., 2022) | Black-box | CVPR-2022 | WGAN-GP | |
| C2FMI (Ye et al., 2023) | Black-box | TDSC-2023 | StyleGAN2 | |
| RLBMI (Han et al., 2023) | Black-box | CVPR-2023 | WGAN-GP | |
| LOKT (Nguyen et al., 2023a) | Black-box | NeurIPS-2023 | ACGAN |
Defenses. To effectively defend against MI attacks, the defender typically employs defense strategies during the training process of victim classifiers. Our benchmark includes 4 typical defense strategies and the details are presented in Table 2.
| Defense Strategy | Reference | Core Technique | Description |
| MID (Wang et al., 2021c) | AAAI-2021 | Regularization | Utilize mutual information regularization to limit leaked information about the model input in the model output |
| BiDO (Peng et al., 2022) | KDD-2022 | Regularization | Minimize dependency between latent vectors and the model input while maximizing dependency between latent vectors and the outputs |
| LS (Struppek et al., 2023) | ICLR-2024 | Label Smoothing | Adjusting the label smoothing with negative factors contributes to increasing privacy protection |
| TL (Ho et al., 2024) | CVPR-2024 | Transfer Learning | Utilize transfer learning to limit the number of layers encoding privacy-sensitive information for robustness to MI attacks |

### v3

# MIBench: A Comprehensive Framework for Benchmarking Model Inversion Attack and Defense

@@ Abstract
Model Inversion (MI) attacks aim at leveraging the output information of target models to reconstruct privacy-sensitive training data, raising critical concerns regarding the privacy vulnerabilities of Deep Neural Networks (DNNs). Unfortunately, in tandem with the rapid evolution of MI attacks, the absence of a comprehensive benchmark with standardized metrics and reproducible implementations has emerged as a formidable challenge. This deficiency has hindered objective comparison of methodological advancements and reliable assessment of defense efficacy. To address this critical gap, we build the first practical benchmark named MIBench for systematic evaluation of model inversion attacks and defenses. This benchmark bases on an extensible and reproducible modular-based toolbox which currently integrates a total of 19 state-of-the-art attack and defense methods and encompasses 9 standardized evaluation protocols. Capitalizing on this foundation, we conduct extensive evaluation from multiple perspectives to holistically compare and analyze various methods across different scenarios, such as the impact of target resolution, model predictive power, defense performance and adversarial robustness.
@@ 1 Introduction
Model Inversion (MI) attacks aim to infer private training data from the output information of target models. In recent years, the proliferation of MI attacks have raised alarms over the potential privacy breaches of sensitive personal information, including the leakage of privacy images in face recognition models [17], sensitive health details in medical data [51], financial information such as transaction records and account balances [36], and personal preferences and social connections in social media data [10]. Since [64] first introduced the Generative Adversarial Networks (GANs) as stronger image priors and established the first GAN-based MI framework, a series of subsequent studies [2, 61, 50, 39] have driven substantial progress in this emerging field during the past few years.
@@
Despite the remarkable development in the MI field, some critical challenges are posed due to the absence of a comprehensive and reliable benchmark. 1) Limited Reproducibility: While most MI methods provide open-source implementations, significant reproducibility challenges persist. First, heterogeneous runtime environments [43, 60, 49] impede simultaneous replication across different studies. Second, the unavailability of pre-trained models in some works [63, 3, 49] hinders comprehensive comparative experimentation. 2) Inadequate and Unfair comparison on different methods: The evaluation of new methods is often confined to comparisons with a narrow selection of prior works, limiting the scope and depth of analysis. Some methods [35, 60] exhibit superior performance for lower-resolution images while other methods [43, 39] perform better at higher resolutions. However, these studies only perform evaluation under their predominant resolutions, and thus do not provide a unified and comprehensive comparison. These shortcomings impede both the accurate measurement of advancements in the MI field and the systematic exploration of its theoretical underpinnings, underscoring the urgent need for a harmonized framework to facilitate robust and transparent research practices.
@@
To alleviate these problems, we establish the first comprehensive benchmark named MIBench for MI attacks and defenses. Our proposed MIBench is built upon an extensible modular-based toolbox, which has encompassed a total of 19 distinct attack and defense methods and coupled with 9 prevalent evaluation protocols to adequately measure the comprehensive performance of individual MI methods. Furthermore, we conduct extensive evaluation from multiple perspectives to achieve a thorough appraisal for existing MI methods, while simultaneously venturing into undiscovered insights to inspire potential avenues for future research. We expect that this reproducible and reliable benchmark will facilitate the further development of MI field and bring more innovative explorations in the subsequent study. Our main contributions are as follows:
We build the first comprehensive benchmark in MI field, featuring a reliable and reproducible modular-based toolbox and comprehensive evaluation on multiple scenarios.
We currently implement a total of 19 state-of-the-art attack methods and defense strategies and 9 standarized evaluation protocols in our toolbox.
We conduct extensive experiments to thoroughly assess different MI attacks and defenses under multiple settings and investigate the effects of different factors to offer new insights on the MI field.
@@ 2 Related Work
Model Inversion Attacks. In MI attacks, the malicious adversary aims to reconstruct privacy-sensitive data by leveraging the output prediction confidence of target classifiers and other necessary auxiliary priors. Normally, the attacker requires a public dataset that shares structural similarities but has no overlapping labels with the private dataset, which is utilized to pre-train a generator before launching an attack. For example, an open-source face dataset serves as essential public data when targeting a face recognition model. In typical MI attacks based on Generative Adversarial Networks (GANs), attackers attempt to recover private images from the GAN’s latent vectors initialized by Gaussian distribution, given the target image classifier parameterized with weights and the trained generator . The attack process is formulated as follows:
@@
where is the target class, typically denotes the classification loss, is a hyperparameter, and is the prior knowledge regularization (e.g., the discriminator’s classification loss) used to improve the image fidelity of . Following the above optimization, the adversary updates the latent vectors into the optimal results and generate final images through .
@@
Based on the accessibility of target models, MI attacks can be further split into white-box attacks, black-box attacks, and label-only attacks. In white-box settings, the full parameters of target models are available to attackers. GMI [63] first proposes to incorporate the rich prior knowledge [8, 15, 9] within the pre-trained GANs [13]. Specifically, GMI starts by generating a series of preliminary fake images, and then iteratively optimizes the input latent vectors that are used for generation. Based on GMI, KEDMI [3] refines the discriminator by introducing target labels to recover the distribution of the input latent vectors. VMI [49] utilizes variational inference to model MI attacks and adopts KL-divergence as the regularization to better estimate the target distribution. PPA [43] introduces a series of techniques such as initial selection, post-selection, and data argumentation to enhance MI attacks and manages to recover high-resolution images by the pre-trained StyleGAN2 [22]. LOMMA [35] integrates model augmentation and model distillation into MI attacks to tackle the problem of over-fitting. PLGMI [60] leverages a top- selection technique to generate pseudo labels to further guide the training process of GAN. The most recent work IFGMI[39] delves into the internal architecture of GANs and achieve higher-quality image inversion through optimization of the intermediate-layer feature vectors.
In black-box settings, the attackers only have access to the output prediction confidence vectors (also named soft-label). Thus, gradients of the target model can no longer be computed by the back-propagation process. Mirror [1] addresses this problem by employing a genetic algorithm as an alternative to gradient descent for optimization. Following [1], C2FMI [59] designs an inverse net to map soft labels into the latent space of pre-trained GANs, which serves as coarse optimization before launching the genetic algorithm. RLBMI [16] presents a novel paradigm by incorporating reinforcement learning as the optimization mechanism.
The label-only settings step further on the black-box settings, relaxing the accessibility to merely one-hot vectors containing the final predicted class (also named hard-label). BREPMI [21] estimates the optimization direction via the boundary repulsion technique. LOKT[34] proposes to conduct MI attacks on various surrogate models instead of the unfamiliar victim model, transforming the label-only settings into white-box settings.
@@
Model Inversion Defenses. To defend MI attacks, most existing methods can be categorized into two types according to their functional stage: train-time defense [12, 47, 29, 52, 37, 44] and inference-time defense [56, 53, 58]. Train-time defense refers to incorporating the defense strategies during the training process. MID [52] proposes penalizing the mutual information between model inputs and outputs in the training loss to reduce the redundant information carried in the model output that may be abused by the attackers. However, simply decreasing the dependency between the inputs and outputs also results in model performance degradation. To strike a better balance between model utility and user privacy, BiDO [37] minimizes the dependency between the inputs and outputs while maximizing the dependency between the latent representations and outputs. [12] propose to leverage GAN to generate fake public samples to mislead the attackers. [47] defend MI attacks by adding Laplacian noise to intermediate representations. LS [44] finds that label smoothing with negative factors can help privacy preservation. TL [20] leverages transfer learning to limit the number of layers encoding sensitive information and thus improves the robustness to MI attacks. The most recent work [25] points out that the skip connection structure of DNNs enhances MI attacks and first proposes defenses from the perspective of robust model architecture.
Inference-time defense refers to reducing the private information carried in the victim model’s output to promote privacy without modifying the training process. This type of defense often functions as an additional plugin. Purifier [57] proposed to train an autoencoder to purify the output vector by increasing its degree of dispersion. SSD [65] proposed to minimize the conditional mutual information under the constraint of bounded distortion.
| GMI [63] | White-box | CVPR-2020 | WGAN | |
| KEDMI [3] | White-box | ICCV-2021 | Inversion-specific GAN† | |
| VMI [49] | White-box | NeurIPS-2021 | StyleGAN2 | |
| PPA [43] | White-box | ICML-2022 | StyleGAN2 | |
| PLGMI [60] | White-box | AAAI-2023 | Conditional GAN | |
| LOMMA [35] | White-box | CVPR-2023 | ||
| IF-GMI [39] | White-box | ECCV-2024 | StyleGAN2 | |
| Mirror∗ [1] | White-box/Black-box | NDSS-2022 | StyleGAN | |
| C2FMI [59] | Black-box | TDSC-2023 | StyleGAN2 | |
| RLBMI [16] | Black-box | CVPR-2023 | WGAN | |
| BREPMI [21] | Label-only | CVPR-2022 | WGAN | |
| LOKT [34] | Label-only | NeurIPS-2023 | ACGAN |
| Defense Strategy | Functional Stage | Reference | Core Technique | Description |
| MID [52] | Train-time | AAAI-2021 | Regularization | Utilize mutual information regularization to limit leaked information about the model input in the model output |
| BiDO [37] | Train-time | KDD-2022 | Regularization | Minimize dependency between latent vectors and the model input while maximizing dependency between latent vectors and the outputs |
| LS [44] | Train-time | ICLR-2024 | Label Smoothing | Adjusting the label smoothing with negative factors contributes to increasing privacy protection |
| TL [20] | Train-time | CVPR-2024 | Transfer Learning | Utilize transfer learning to limit the number of layers encoding privacy-sensitive information for robustness to MI attacks |
| RoLSS/SSF/TTS∗ [25] | Train-time | ECCV-2024 | Skip Connection | Properly remove some skip connection structures in the target classifier during the training stage |
| Purifier [57] | Inference-time | AAAI-2023 | Conditional VAE | Reform confidences and swap labels to make members behave like non-members |
| SSD [65] | Inference-time | ICLR-2025 | Conditional Mutual Information | Minimize conditional mutual information to make predictions less dependent on inputs and more dependent on ground truths |
∗The three defenses proposed by [25] have minor distinction between each other, therefore we select the TTS as the
representative method in our evaluation.
@@ 3.1 Dataset
Considering existing MI attacks primarily focus on reconstructing private facial data from image classifiers, we select 4 widely recognized face datasets as the basic datasets, including Flickr-Faces-HQ (FFHQ) [22], MetFaces [23], FaceScrub [33], and CelebFaces Attributes (CelebA) [30]. Generally, FFHQ and MetFaces are employed as public datasets for pre-training auxiliary priors, whereas FaceScrub and CelebA serve as the target private datasets to attack. Our benchmark is modularized to facilitate researchers to freely combine public datasets with private ones, enabling customized experimental setups. Extensive evaluation on more non-facial datasets (Stanford Dogs [5] and AFHQ-Dog [4]) are presented in Supp.C.3.
Notably, the target resolutions across different MI attacks are not uniform. The majority of attack methods concentrate on low-resolution images of , while recent attack methods have begun to focus on higher resolutions, such as . Therefore, our benchmark offers 2 versions of low-resolution and higher-resolution for the aforementioned 4 datasets and prepares multiple transformation tools for processing images, freeing researchers from the laborious tasks of data preprocessing. More details regarding the datasets can be found in the Supp.B.2.
@@ 3.2 Implemented Methods
Our benchmark includes a total of 19 methods, comprising 12 attack methods and 7 defense strategies. With a focus on Generative Adversarial Network (GAN)-based MI attacks, we selectively reproduce methods from recent years that have been published in top-tier conferences or journals in the computer vision or machine learning domains. This criterion ensures the reliability and validity of the implemented methods. Considering the main targets in our benchmark are image classifiers for RGB images, the learning-based MI attacks [11, 42, 55] are not incorporated currently. More detailed information about the implemented methods is stated in Supp.A.2 and Supp.A.3.
@@
Defenses. To effectively defend against MI attacks, the defender typically employs defense strategies during the training process of victim classifiers. Our benchmark has implemented 4 typical defense strategies and corresponding details are presented in Table 2.